OpenStudy (anonymous):

f(x) = x3 + 3x2 - x - 3 Part A: What are the factors of f(x)? Show your work. (3 points) Part B: What are the zeros of f(x)? Show your work. (2 points) Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x). (5 points) a few moments ago

OpenStudy (shamil98):

By looking at the highest exponent in your function you can determine that you will have three multiples.. (x )(x )(x ) use synthetic division , the numbers that end up with zero as the remainder are your factors.

OpenStudy (math&ing001):

A: f(x)=x2 (x+3) - (x+3) factor by (x+3) B: when f(x)= (x-a)(x-b)(x-c).... then a,b, and c are the zeros